# Generalized Einstein-Rosen bridge inside black holes

### 26 Dec 2023 v1

@@ Abstract
We generalize the notion of Einstein-Rosen bridge by defining it as a space-ilke connection between two universes with regions of asymptotically minkowskian space-time infinities. The corresponding symmetry and asymmetry properties of the generalized Einstein-Rosen bridge are considered at the cases of Reissner-Nordström and Kerr metrics. We elucidate the versatility of intriguing symmetry and asymmetry phenomena outside and inside black holes. For description of the test particle (planet and photon) motion it is used the Kerr-Newman metric of the rotating and electrically charged black hole. In particular, it is demonstrated the symmetry and asymmetry of the one-way Einstein-Rosen bridge inside black hole toward and through the plethora of endless and infinite universes.
@@ 1 Introduction
In this paper we generalize the notion of Einstein-Rosen bridge by defining it as a space-ilke connection between two universes with regions of asymptotically minkowskian space-time infinities. The corresponding symmetry and asymmetry properties of the generalized Einstein-Rosen bridge are considered at the cases of Reissner-Nordström and Kerr metrics. We elucidate the versatility of intriguing symmetry and asymmetry phenomena outside and inside black holes. For description of the test particle (planet and photon) motion it is used the Kerr-Newman metric of the rotating and electrically charged black hole. In particular, it is demonstrated the symmetry and asymmetry of the one-way Einstein-Rosen bridge inside black hole toward and through the plethora of endless and infinite universes.
@@
Figure 1: Carter–Penrose diagram for the spherically symmetric Reissner–Nordström black hole with electric charge $q=0.99$. The spaceship starts from the point $A$ at $R_{+}$-region toward its multi-planetary future inside the black hole. The astronauts are planning to use the Einstein-Rosen bridge magenta curve) and intersect both the black hole event horizon $r_{+}$ and Cauchy horizon $r_{-}$ at finite their proper time. After appearing near the black hole singularity at $r=0$, the space ship uses its powerful engines to change the direction of motion and escape the tidal destruction at small radii. In result, the voyage is happily finishing at point $B$ (may be at the Earth-like planet) in other infinite universe. The symmetry is in possibility to repeat the complete route of this voyage staring from the point $B$ but only in the forward direction in time toward the another multi-planetary future. It is impossible to return the native Earth due to impossibility of any motion beyond the light cone. This is the motion asymmetry on the one-way Einstein-Rosen bridge inside black hole.
@@
Figure 2: $2D$ presentation of the voyage through the Reissner–Nordström black hole interiors by using the Einstein-Rosen bridge. This picture is geometrically absolutely symmetric or, in other words, it is nicely symmetric. At the same time, this picture is misleading and physically controversial: Indeed. the voyage is starting at apogee $r_{a}$ from the position at point ${\color[rgb]{1,0,1}A}$, then reach the perigee $r_{p}$ and return the apogee $r_{a}$ at the point $B$ for a finite proper time, demonstrating the absolute geometric symmetry. Meanwhile, there is a crucial hitch: this apogee $r_{a}$ at the point ${\color[rgb]{1,0,1}B}$ is not in the native universe, but in the other quite distant universe, as it is clearly viewed at the Carter–Penrouse diagram at Fig. 1. The apogee $r_{a}$ and perigee $r_{p}$ radii are shown by dashed circles. Respectively, the event radii of event horizon $r_{+}$ and Cauchy horizon $r_{-}$ are shown by solid circles. The magenta curve here and at the Fig. 4 is numerically calculated geodesic trajectory with using equations of motion (3)–(6) for massive test particles ($\mu\neq 0$).
@@
Figure 5: Trajectory of the test planet with parameters $\gamma=0.85$, $\lambda=1.7$ and $Q=1$, which is plunging into the fast-rotating Kerr black hole with spin $a=0.9982$. This planet is winding up on the black hole event horizon higher the equatorial plane. The blue curve here and in the Fig. 6 is the numerically calculated geodesic trajectory with using equations of motion (3)–(6) for massless test particles like photons ($\mu=0$).
Figure 6: Trajectory of the test planet with parameters $\gamma=0.85$, $\lambda=1.7$ and $Q=1$, which is plunging into the fast-rotating Kerr black hole with spin $a=0.9982$. This planet is winding up on the black hole event horizon higher the equatorial plane. The blue curve here and in the Fig. 7 is the numerically calculated geodesic trajectory with using equations of motion (3)–(6) for massless test particles like photons ($\mu=0$).
Figure 7: Trajectory of the photon with parameters $\lambda=2$ and $Q=1$, which is plunging into the fast-rotating Kerr black hole with $a=0.9982$ and is winding up on the black hole event horizon below the equatorial plane. The multicolored curve is the numerically calculated geodesic trajectory with using equations of motion (3)–(6) for massless test particles like photons ($\mu=0$).
@@ 2 Kerr-Newman metric for rotating and electrically charged black hole
$M$ — black hole mass, $q$ — black hole electric charge, $a=J/M$ — specific black hole angular momentum (spin). The two roots of equation $\Delta=0$ are $r_{+}=1+\sqrt{M^{2}-a^{2}-q^{2}}$ — the event horizon of the black hole and $r_{-}=1-\sqrt{M^{2}-a^{2}-q^{2}}$ — the Cauchy horizon of the black hole.
@@
In the following Sections we will describe the symmetry and asymmetry of test object motion in the gravitational field of the Kerr-Newman black hole. We use equations of motion in the Kerr–Newman metric (3)–(6) in our analytic and numerical calculations of test particle geodesic trajectories [13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
@@ 3 One-way Einstein-Rosen bridge inside black hole
$2D$ presentation of the voyage through the interiors of Reissner–Nordström black hole with interiors by using the Einstein-Rosen bridge is shown at Fig. 2. The electric charge of the black hole is $e=0.99$. A test planet (or spaceship) with the electric charge $\varepsilon=-1.5$ is periodically orbiting around black holes with orbital parameters $\gamma=0.5$, $\lambda=0.5$, corresponding to the maximal radius (apogee) $r_{\rm max}=1.65$ and minimal radius (perigee) $r_{\rm min}=0.29$, respectively, in dimensional units. The periodic planet geodesic trajectories (magenta curves, both at Fig. 2 and Fig. 4), were calculated numerically with using equations of motion (3)–(6) for massive test particles ($\mu\neq 0$). Note, that the periodical motion of the test planet is limited in time due to energy losses in inevitable emission of the gravitational waves.
The picture at Fig. 2 is geometrically absolutely symmetric or, in other words, it is completely or nicely symmetric. The geodesic trajectories of test planet ($\mu\neq 0$) at this Fig. 2 and at Fig. 3 (the red-colored curves), are numerically calculated [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], by using the corresponding equations of motion in the Kerr-Newman metric (3)–(6).
@@ 4 Symmetry and asymmetry of test particle trajectories near rotating black hole
Figs. 57 demonstrate both symmetry and asymmetry features of massive and massless particle trajectories plunging into rotating Kerr black hole with spin $a=0.9982$. Magenta arrows shows the direction of the black hole rotation in accordance with the gimlet rule. The multicolored curves at Figs. 7 and 8 are the geodesic trajectories for massless test particles like photons ($\mu=0$) numerically calculated with using equations of motion (3)–(6). By approaching the black hole, the trajectories of all particles, both massive and massless ones, are infinitely winding up on the black hole event horizon in the direction of the black hole rotation and at the fixed latitudes. This winding up is a manifestation of symmetry of behavior of all trajectories, plunging into rotating black hole. At the same the direction of the black hole rotation is a corresponding manifestation of asymmetry of the gravitational field of the Kerr metric.
At last, for completeness of black hole symmetric and asymmetric properties, at Figs. 8 is shown the trajectory of the test planet ($\mu\neq 0$) with parameters $\gamma=1$, $\lambda=1$ and $Q=0.5$. This test planet is plunging into the spherically symmetric and nonrotating Schwarzschild black hole (with both the spin $a=0$ and electric charge $q=0$), starting from the radial distance $r=6$. It must be especially checked that the traversable (though only one-way in time and direction) Einstein-Rosen bridge is absent at all inside the Schwarzschild black hole (see for details, e.  g., [4, 6]).
@@ 5 Conclusion and Discussion
It is demonstrated the symmetry and asymmetry of the voyage on one-way Einstein-Rosen bridge inside black hole toward the endless multiplanetary future. The apparent symmetry of both the Carter–Penrose and embedding diagrams is mainly related with a pure geometrical vision of this phenomenon. Quite the contrary, the physical (space-time) vision elucidates the absolute asymmetry of the Einstein-Rosen bridge due to existence of the light-cone limitation for possible motions.
Note, that the traversable (though only one-way in time and direction) Einstein-Rosen bridge exist only in the case of both rotating Kerr $a\neq 0$ and electrically charged Reissner–Nordström $q\neq 0$ black holes. It is absent at all inside the Schwarzschild black hole (see for details, e.  g., [4, 6])
The infinite winding up of trajectories of all particles on the black hole event horizon is a manifestation of symmetry of behavior of all trajectories, plunging into rotating black hole. At the same the direction of the black hole rotation is a corresponding manifestation of asymmetry of the gravitational field of the Kerr metric.
We are grateful to E. O. Babichev, V. A. Berezin, Yu. N. Eroshenko and A. L. Smirnov for stimulating discussions.

### 15 Jan 2024 v2

@@ Abstract
We generalize the notion of Einstein-Rosen bridge by defining it as a space-like connection between two universes with regions of asymptotically minkowskian space-time infinities. The corresponding symmetry and asymmetry properties of the generalized Einstein-Rosen bridge are considered at the cases of Reissner-Nordström and Kerr metrics. We elucidate the versatility of intriguing symmetry and asymmetry phenomena outside and inside black holes. For description of the test particle (planet and photon) motion it is used the Kerr-Newman metric of the rotating and electrically charged black hole. In particular, it is demonstrated the symmetry and asymmetry of the one-way Einstein-Rosen bridge inside black hole toward and through the plethora of endless and infinite universes.
@@
Keywords: gravitation, General Relativity, black holes
@@ 1 Introduction
In this paper we generalize the notion of Einstein-Rosen bridge by defining it as a space-ilke connection between two universes with regions of asymptotically minkowskian space-time infinities. The corresponding symmetry and asymmetry properties of the generalized Einstein-Rosen bridge are considered at the cases of Reissner-Nordström and Kerr metrics. We elucidate the versatility of intriguing symmetry and asymmetry phenomena outside and inside black holes. For description of the test particle (planet and photon) motion it is used the Kerr-Newman metric of the rotating and electrically charged black hole. It is demonstrated the symmetry and asymmetry of the one-way Einstein-Rosen bridge inside black hole toward and through the plethora of endless and infinite universes.
@@
Figure 1: Carter–Penrose diagram for the spherically symmetric Reissner–Nordström black hole with electric charge $q=0.99$. The spaceship starts from the point ${\color[rgb]{1,0,1}A}$ at $R_{+}$-region toward its multi-planetary future inside the black hole. Astronauts are planning to use the Einstein-Rosen bridge magenta curve) and intersect both the black hole event horizon $r_{+}$ and Cauchy horizon $r_{-}$ at finite their proper time. After appearing near the black hole singularity at $r=0$, the spaceship uses its powerful engines to change the direction of motion and escape the tidal destruction at small radii. In result, the voyage is happily finishing at point $B$ (may be at the Earth-like planet) in another infinite universe. The symmetry is in possibility to repeat the complete route of this voyage staring from the point ${\color[rgb]{1,0,1}B}$ but only in the forward direction in time toward another multi-planetary future. It is impossible to return the native Earth due to impossibility of any motion beyond the light cone. This is the motion asymmetry on one-way Einstein-Rosen bridge inside black hole.
@@
Figure 2: $2D$ presentation of the voyage through the Reissner–Nordström black hole interiors by using the Einstein-Rosen bridge. This picture is geometrically absolutely symmetric or, in other words, it is nicely symmetric. At the same time, this picture is misleading and physically controversial: Indeed. the voyage is starting at apogee $r_{a}$ from the position at point ${\color[rgb]{1,0,1}A}$, then reach the perigee $r_{p}$ and return the apogee $r_{a}$ at the point $B$ for a finite proper time, demonstrating the absolute geometric symmetry. Meanwhile, there is a crucial hitch: this apogee $r_{a}$ at the point ${\color[rgb]{1,0,1}B}$ is not in the native universe, but in the other quite distant universe, as it is clearly viewed at the Carter–Penrose diagram at Fig. 1. The apogee $r_{a}$ and perigee $r_{p}$ radii are shown by dashed circles. Respectively, the event radii of event horizon $r_{+}$ and Cauchy horizon $r_{-}$ are shown by solid circles. The magenta curve here and at the Fig. 4 is numerically calculated geodesic trajectory with using equations of motion (3)–(6) for massive test particles ($\mu\neq 0$).
@@
Figure 5: Trajectory of the test planet with parameters $\gamma=0.85$, $\lambda=1.7$ and $Q=1$, which is plunging into the fast-rotating Kerr black hole with spin $a=0.9982$. This planet is winding up on the black hole event horizon higher the equatorial plane. Blue curve here and in the Fig. 6 is the numerically calculated geodesic trajectory with using equations of motion (3)–(6) for massless test particles like photons ($\mu=0$).
Figure 6: Trajectory of the test planet with parameters $\gamma=0.85$, $\lambda=1.7$ and $Q=1$, which is plunging into the fast-rotating Kerr black hole with spin $a=0.9982$. This planet is winding up on the black hole event horizon higher the equatorial plane. Blue curve here and in the Fig. 7 is the numerically calculated geodesic trajectory with using equations of motion (3)–(6) for massless test particles like photons ($\mu=0$).
Figure 7: Trajectory of the photon with parameters $\lambda=2$ and $Q=1$, which is plunging into the fast-rotating Kerr black hole with $a=0.9982$ and is winding up on the black hole event horizon below the equatorial plane. Multicolored curve is the numerically calculated geodesic trajectory with using equations of motion (3)–(6) for massless test particles like photons ($\mu=0$).
@@ 2 Kerr-Newman metric for rotating and electrically charged black hole
$M$  —  black hole mass, $q$  —  black hole electric charge, $a=J/M$  —  specific black hole angular momentum (spin). The two roots of equation $\Delta=0$ are  $r_{+}=1+\sqrt{M^{2}-a^{2}-q^{2}}$  —  the black hole event horizon and $r_{-}=1-\sqrt{M^{2}-a^{2}-q^{2}}$  —  the black hole Cauchy horizon.
@@
In the following Sections we will describe the symmetry and asymmetry of test object motion in the gravitational field of the Kerr-Newman black hole. We use equations of motion in the Kerr–Newman metric (3)–(6) in our analytic and numerical calculations of test particle geodesic trajectories [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
@@ 3 One-way Einstein-Rosen bridge inside black hole
$2D$ presentation of the voyage through interiors of Reissner–Nordström black hole by using the Einstein-Rosen bridge is shown at Fig. 2. The electric charge of the black hole is $e=0.99$. A test planet (or spaceship) with the electric charge $\varepsilon=-1.5$ is periodically orbiting around black holes with orbital parameters $\gamma=0.5$, $\lambda=0.5$, corresponding to the maximal radius (apogee) $r_{\rm max}=1.65$ and minimal radius (perigee) $r_{\rm min}=0.29$, respectively, in dimensional units. The periodic planet geodesic trajectories (magenta curves, both at Fig. 2 and Fig. 4), were calculated numerically with using equations of motion (3)–(6) for massive test particles ($\mu\neq 0$). Note, that the periodical motion of the test planet is limited in time due to energy losses in inevitable emission of the gravitational waves.
The picture at Fig. 2 is geometrically absolutely symmetric or, in other words, it is completely or nicely symmetric. The geodesic trajectories of test planet ($\mu\neq 0$) at this Fig. 2 and at Fig. 3 (the red-colored curves), are numerically calculated [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], by using the corresponding equations of motion in the Kerr-Newman metric (3)–(6).
@@ 4 Symmetry and asymmetry of test particle trajectories near rotating black hole
Figs. 57 demonstrate both symmetry and asymmetry features of massive and massless particle trajectories plunging into rotating Kerr black hole with spin $a=0.9982$. Magenta arrows show the direction of the black hole rotation in accordance with the gimlet rule. Multicolored curves at Figs. 7 and 8 are the geodesic trajectories for massless test particles like photons ($\mu=0$) numerically calculated with using equations of motion (3)–(6). By approaching the black hole, the trajectories of all particles, both massive and massless ones, are infinitely winding up on the black hole event horizon in the direction of black hole rotation and at the fixed latitudes. This winding up is a symmetry manifestation of all trajectories behavior, which are plunging into rotating black hole. At the same direction of black hole rotation is a corresponding asymmetry manifestation of the gravitational field of the Kerr metric.
At last, for completeness of black hole symmetric and asymmetric properties, at Figs. 8 is shown the trajectory of the test planet ($\mu\neq 0$) with parameters $\gamma=1$, $\lambda=1$ and $Q=0.5$. This test planet is plunging into the spherically symmetric and nonrotating Schwarzschild black hole (with both the spin $a=0$ and electric charge $q=0$), starting from the radial distance $r=6$.
It must be especially checked that the traversable (though only one-way in time and direction) Einstein-Rosen bridge is absent at all inside the Schwarzschild black hole (see for details, e.  g., [4, 6]).
@@ 5 Conclusion and Discussion
It is demonstrated the symmetry and asymmetry of the voyage on one-way Einstein-Rosen bridge inside black hole toward the endless multiplanetary future.
The apparent symmetry of both the Carter–Penrose and embedding diagrams is mainly related with a pure geometrical vision of this phenomenon. Quite the contrary, the physical (space-time) vision elucidates the absolute asymmetry of the Einstein-Rosen bridge due to existence of the light-cone limitation for possible motions.
Note, that the traversable (though only one-way in time and direction) Einstein-Rosen bridge exist only in the case of both rotating Kerr $a\neq 0$ and electrically charged Reissner–Nordström $q\neq 0$ black holes. It is absent at all inside the Schwarzschild black hole (see for details, e.  g., [4, 6]).
The infinite winding up of particle trajectories on the black hole event horizon is a manifestation of symmetry behavior of all trajectories, plunging into rotating black hole. At the same direction of the black hole rotation is a corresponding manifestation of asymmetry of the gravitational field of the Kerr metric.
We are grateful to E. O. Babichev, V. A. Berezin, Yu. N. Eroshenko, N. O. Nazarova and A. L. Smirnov for stimulating discussions.